OpenStudy (anonymous):

What is the exact value of the expression √12 + √50 - √48 ? Simplify completely. √14 5√2 - 2√3 25√2 - 12√3 √50 - √12

OpenStudy (poohbear):

I got 2√3 + 5√2 - 4√3. .

OpenStudy (anonymous):

but which one

OpenStudy (poohbear):

Oh, I didn't know those were the choices. :P But it should be be 5√2 - 2√3.

OpenStudy (anonymous):

yea, that's what i got

OpenStudy (anonymous):

ok thanks

OpenStudy (radar):

\[\sqrt{12}=\sqrt{4}\sqrt{3}=2\sqrt{3}\]\[\sqrt{50}=\sqrt{25}\sqrt{2}=5\sqrt{2}\]\[\- sqrt{48}=-\sqrt{16}\sqrt{3}=-4\sqrt{3}\]

OpenStudy (radar):

Combining similar radicands getting:\[5\sqrt{2}-2\sqrt{3}\]